OpenStudy (anonymous):

a machine needs 2 of its gears replaced . you find 2 gears one manufactured 5 years ago and the other manufactured 10 years ago. the specifications sheet says that the chance of failure is 5 % after 5 years and increases to 8% after 10 years . if the gears on the machine are in no way linked to each other what are the chances that both will fail?

OpenStudy (phi):

you can multiply the chance of failure

OpenStudy (anonymous):

so iit would be 0.4% or 40%

OpenStudy (phi):

chance of failure for one of the gears is 5% which is the same as 5/100 or 0.05 the other is 8% = 8/100 = 0.08 multiply those

OpenStudy (anonymous):

szo iitsz 0.4 thanx ii hope iim riight lol

OpenStudy (phi):

it depends if you write the answer as a percent or as a decimal. which way do they want it ?

OpenStudy (anonymous):

%

OpenStudy (phi):

the easiest way to not get confused is change everything to decimals and mulitply 0.05*0.08= 0.0040 then change back to a percent by multiplying by 100 (and adding a % sign) 100* 0.0040= 0.4 so we get 0.4 % chance that both gears will fail